# Borsuk-Ulam and Replicable Learning of Large-Margin Halfspaces

### v3

@@ 1.1 Main contributions
For any fixed margin $\gamma\in(0,1)$, finite dimension $d>1$, and accuracy parameter $\epsilon\in(0,1/2)$,
@@
In particular, $\mathcal{H}_{\gamma}^{\infty}$ is not list replicable (equiv. globally stable).
Theorem 1.6 reveals a surprising distinction in the partial setting: list replicability does not follow from bounded Littlestone dimension, replicability, DP-learnability, or even pure DP-learnability. Specifically, for every $\gamma\in(0,1)$, the partial concept class $\mathcal{H}^{\infty}_{\gamma}$ is shared-randomness replicable, pure DP-learnable, and has a finite Littlestone dimension, but it is not list replicable.
The lower bound in Theorem 1.6 relies on a topological argument involving covers of the sphere by antipodal-free open sets. In particular, we apply the local Borsuk-Ulam theorem of [CCMY24], which states that in such a cover, there is a point that belongs to at least $\frac{d}{2}+1$ sets. Alternatively, one could use Ky Fan’s classical theorem [Fan52], but this would yield the slightly weaker bound of $\frac{d}{2}$.
@@
Disambiguations of Large-margin Half-spaces:
One of the central questions studied in [AHHM22] is whether the learnability of a partial concept class can always be inferred from the learnability of a suitable extension of it to a total concept class.
Formally, a disambiguation of a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is a total concept class ¯{±$\overline{\mathcal{H}}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$ such that for every $h\in\mathcal{H}$ and every finite $S\subseteq h^{-1}(\left\{\pm 1\right\})$, there exists an $\bar{h}\in\overline{\mathcal{H}}$ that is consistent with $h$ on $S$. Intuitively, one may think of disambiguation as replacing each $\star$ with $-1$ or $+1$, although this intuition is not completely rigorous in the infinite case.
Note that any distribution realizable by a partial concept class $\mathcal{H}$ is also realizable by any disambiguation $\overline{\mathcal{H}}$ of $\mathcal{H}$. Consequently, if a disambiguation $\overline{\mathcal{H}}$ is efficiently learnable, the same holds true for $\mathcal{H}$.
Motivated by the fact that $\operatorname{Ldim}(\mathcal{H}_{\gamma}^{d})\leq\gamma^{-2}$, [AHHM22] posed the following question.
Question 1.7 ([AHHM22, Question 4]).
Does there exist a disambiguation of $\mathcal{H}_{\gamma}^{d}$ by a total concept class whose VC/Littlestone dimension is bounded by a function of $\gamma$?
In the following theorem, we provide a negative answer to the Littlestone dimension part of 1.7.
@@
This follows by combining the lower bound of Theorem 1.6 and a result of [GGKM21] showing that for every total concept class $\mathcal{H}$, $\operatorname{LR}_{\epsilon}(\mathcal{H})\leq 2^{O_{\epsilon}\left(\operatorname{Ldim}(\mathcal{H})^{2}\right)}$. ∎
On the other hand, there is a disambiguation $\overline{\mathcal{H}}$ of $\mathcal{H}_{\gamma}^{d}$ with Littlestone dimension $\operatorname{Ldim}(\overline{\mathcal{H}})=O(d)$ since [HHM23] shows that there is a disambiguation of $\mathcal{H}_{\gamma}^{d}$ into a finite total concept class $\overline{\mathcal{H}}$ with $2^{O(d)}$ concepts. The upper bound on the Littlestone dimension follows from the fact that every finite class $\mathcal{H}$ satisfies $\operatorname{VCdim}(\mathcal{H})\leq\operatorname{Ldim}(\mathcal{H})\leq\log|\mathcal{H}|$.
Theorem 1.8 has two important implications. Firstly, it shows $\mathcal{H}^{\infty}_{\gamma}$ is a partial class with finite Littlestone dimension, but any disambiguation of it has an infinite Littlestone dimension. The existence of such partial classes was posed as a question in [AHHM22] and answered in [CHHH23] through a complex construction based on Göös’ breakthrough [Göö15] in communication complexity. Theorem 1.8 offers a more natural example of a concept class with this property.
Secondly, the construction of  [CHHH23] has Littlestone dimension $2$, and in [CHHH23, Question 4.1], they asked whether there exists a partial concept class with Littlestone dimension $1$ such that all its disambiguations have infinite Littlestone dimension. The following corollary of Theorem 1.8 resolves their question.
@@
Let $\gamma\in(1/\sqrt{2},1)$. Then $\operatorname{Ldim}(\mathcal{H}_{\gamma}^{\infty})=1$, while any disambiguation of $\mathcal{H}_{\gamma}^{\infty}$ has infinite Littlestone dimension.
@@
We have $21\leq\operatorname{Ldim}(\mathcal{H}_{\gamma}^{\infty})\leq 1/\gamma^{2}<$ 2. An application of Theorem 1.8 completes the proof. ∎
List replicability of finite classes.
In [CMY23, Theorems 5 and 13], Chase, Moran, and Yehudayoff proved that the list replicability number of any finite set of homogeneous half-spaces and points that do not lie on the half-spaces in $\mathbb{R}^{2}$ is at most $2$. They further asked whether a similar finite bound holds in higher dimensions. As we discuss below, we give an affirmative answer to their question in Theorem 1.11.
Definition 1.10.
@@
Define the finitary list replicability number of a concept class {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$, denoted $\widetilde{\operatorname{LR}}(\mathcal{H})$, as
$\widetilde{\operatorname{LR}}(\mathcal{H})\coloneqq\sup_{\text{finite }S\subseteq\mathcal{X}}\operatorname{LR}(\mathcal{H}|_{S}).$
Given $d\in\mathbb{N}$, let $\mathcal{H}^{d}$ be the partial concept class defined by points in $\mathbb{S}^{d-1}$ and partial concepts $h_{w}:\mathbb{S}^{d-1}\to\left\{\pm 1,\star\right\}$, where $w\in\mathbb{S}^{d-1}$, given by
$h_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}(\langle w,x\rangle)&\text{if }\langle w,x\rangle\neq 0\\ \star&\text{otherwise}\end{cases}.$
Note that for every $\gamma>0$, $\mathcal{H}^{d}_{\gamma}$ is a subclass of $\mathcal{H}^{d}$.
Since $\mathcal{H}^{2}$ has infinite threshold dimension [AHHM22, Definition 28], by the celebrated result of [ABL+22], it is not globally stable and therefore $\operatorname{LR}(\mathcal{H}^{2})=\infty$. Consequently, the aforementioned result of [CMY23] implies the following interesting gap:
@@
The authors of [CMY23] ask whether a similar bound holds for $\widetilde{\operatorname{LR}}(\mathcal{H}^{3})$. In the following theorem, we answer their question by extending their result to all dimensions.
@@
The upper bound is an easy consequence of the upper bound of our main theorem (Theorem 1.6). Indeed, any finite set of points $S\subset\mathbb{S}^{d-1}$ and hypotheses $H\subset\mathcal{H}^{d}|_{S}$ defined by unit vectors $W\subset\mathbb{S}^{d-1}$ is a sub-concept class of $\mathcal{H}^{d}_{\gamma}$, where $\gamma=\min_{x\in S,w\in W}|\langle w,x\rangle|$. The claim now follows from our upper bound from Theorem 1.6.
For the lower bound, we use a result of Chase, Moran, and Yehudayoff [CMY23, Theorem 3] stating that for every concept class $\mathcal{H}$,
$\operatorname{LR}(\mathcal{H})\geq\operatorname{VCdim}(\mathcal{H}).$
The result now follows as $\operatorname{VCdim}(\mathcal{H}^{d})=d$, and therefore, $\mathcal{H}^{d}$ has a finite subclass of VC dimension $d$. ∎
The sign-rank of a sign matrix $A$, denoted by ${\mathbf{rk}}_{\pm}(A)$, is the smallest rank of a real matrix $B_{\mathcal{X}\times\mathcal{Y}}$ such that the entries of $B$ are nonzero and have the same signs as their corresponding entries in $A$. The definition of sign-rank naturally extends to partial matrices, where for invalid entries of $A$, the corresponding entry in $B$ could be any real number.
Geometrically, sign-rank is the smallest dimension in which the matrix is realized as points and homogeneous half-spaces. We will state the definition of sign-rank in this terminology for partial classes.
@@
The sign-rank of a partial class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, denoted by ${\mathbf{rk}}_{\pm}(\mathcal{H})$, is the smallest $d$ such that there exist vectors $u_{h},v_{x}\in\mathbb{R}^{d}$ for all pairs $h\in\mathcal{H},x\in\mathcal{X}$ such that $h(x)=\operatorname{sgn}({\left\langle u_{h},v_{x}\right\rangle})$ whenever $h(x)\neq\star$.
Theorem 1.11 and the VC lower bound of [CMY23, Theorem 3] immediately imply the following general bounds on $\widetilde{\operatorname{LR}}(\mathcal{H})$.
@@
For every partial class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, we have
$\operatorname{VCdim}(\mathcal{H})\leq\widetilde{\operatorname{LR}}(\mathcal{H})\leq\operatorname{rank_{\pm}}(\mathcal{H}).$
Gap Hamming Distance.
$\operatorname{GHD}_{\gamma}^{n}(x,y)\coloneqq\begin{cases}\operatorname{sgn}({\left\langle x,y\right\rangle})&\text{if $|{\left\langle x,y\right\rangle}|\geq\gamma n$}\\ \star&\text{otherwise}\end{cases}.$
For fixed $\gamma\in(0,1)$, the public-coin randomized communication complexity of $\operatorname{GHD}_{\gamma}$ is at most a fixed constant depending only on $\gamma$. This places $\operatorname{GHD}_{\gamma}$ in the complexity class $\sf{BPP}_{0}$, which consists of communication problems with constant-cost public-coin randomized protocols. Moreover, $\operatorname{GHD}_{\gamma}$ exhibits certain properties that separate $\sf{BPP}_{0}$ from some complexity classes (e.g., $\sf{UPP}_{0}$) and measures [HHH23, CLV19, HHM23, FGHH25, Son14].
A key caveat of these separations is that $\operatorname{GHD}_{\gamma}$ is a partial function, and whether it extends to any family of total functions with similar properties remained an open problem before this work. One suggested approach [FGHH25] to resolving these questions is to find a disambiguation of $\operatorname{GHD}_{\gamma}$ to a total function with constant public-coin randomized communication complexity. In Corollary 1.15 below, we prove this is impossible: every disambiguation of $\operatorname{GHD}_{\gamma}^{n}$ has public-coin randomized communication complexity at least $\Omega(\log\log n)$.
First, we show that every such disambiguation has an unbounded Littlestone dimension.
Theorem 1.14.
$\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n}).$
We obtain the following corollary by combining Theorem 1.14 with some known results.
Corollary 1.15.
Let $\gamma\in(0,1)$ be a margin parameter. Every family of disambiguation of the Gap Hamming Distance matrices $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$ has public-coin randomized communication complexity $\Omega(\log\log n)$.
The bound follows by combining Theorem 1.14 with the known relationship between Littlestone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity. See Section 4 for the proof.

### v4

@@ 2 Main theorem
For any fixed margin $\gamma\in(0,1)$, dimension $d>1$, and accuracy parameter $\epsilon\in(0,1/2)$,
@@
Hence, $\tfrac{d}{2}+1\leq\operatorname{LR}(\mathcal{H}_{\gamma}^{d})\leq d$.
The lower bound in Theorem 2.1 relies on a topological argument involving covers of the sphere by antipodal-free open sets. In particular, we apply the local Borsuk-Ulam theorem of [CCMY24], which states that in such a cover, there is a point that belongs to at least $\frac{d}{2}+1$ sets. Alternatively, one could use Ky Fan’s classical theorem [Fan52], but this would yield the slightly weaker lower bound of $\frac{d}{2}$.
@@
2.1 Applications
Separation.
In addition to list replicability, differentially private learnability and shared-randomness replicability [^2] In prior works, shared-randomness replicability is referred to simply as replicability. Since we work with multiple replicability notions, we adopt this terminology to emphasize that different executions of the algorithm use the same random seed. are two other well-studied notions of stability in learning theory (see [MSS23] for an overview). For completeness, formal definitions of these concepts appear in the appendix, though we do not rely on them directly in this paper.
Recent advances in learning theory [ALMM19, BLM20, ABL+22, CMY23, ILPS22], sparked by the influential works on differential privacy in PAC learning, have established that for total concept classes, all of these notions coincide and are characterized by the finiteness of the Littlestone dimension. Specifically, for every total concept class $\mathcal{C}\subseteq\{\pm 1\}^{\mathcal{X}}$, the following are equivalent:
  • $\operatorname{Ldim}(\mathcal{C})<\infty$;
  • $\mathcal{C}$ is list replicable;
  • $\mathcal{C}$ is shared-randomness replicable;
  • $\mathcal{C}$ is approximately differentially private (DP)-learnable.
This naturally raises the question of whether these equivalences extend to partial concept classes. The case of large-margin half-spaces has been studied extensively [BDMN05, LNUZ20, BMNS19, KMST20, BCS20, BMS22a, BMS22b, ILPS22, KKL+24], and the following are known:
  • $\operatorname{Ldim}(\mathcal{H}^{d}_{\gamma})<\gamma^{-2}$;
  • $\mathcal{H}^{d}_{\gamma}$ is (pure) DP-learnable with dimension-independent sample complexity;
  • $\mathcal{H}^{d}_{\gamma}$ is shared-randomness replicable with dimension-independent sample complexity.
Nevertheless, Theorem 2.1 shows that despite these strong positive results, the list replicability number of $\mathcal{H}^{d}_{\gamma}$ necessarily grows with $d$. Consequently, we obtain a sharp separation from the total setting: for partial classes, list replicability does not follow from bounded Littlestone dimension, replicability, DP-learnability, or even pure DP-learnability.
Corollary 2.2 (Separation).
There exists a partial concept class $\mathcal{C}$ that is (pure) DP-learnable, shared-randomness replicable, and satisfies $\operatorname{Ldim}(\mathcal{C})<\infty$, yet it is not list replicable.
Proof.
Fix $\gamma\in(0,1)$, and define the class $\mathcal{H}^{\infty}_{\gamma}$ as follows. Each hypothesis in $\mathcal{H}^{\infty}_{\gamma}$ is specified by a unit vector $w$ of arbitrary finite dimension, i.e., $w\in\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}$. For $x\in\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}$, define
$c_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}({\left\langle w,x\right\rangle})&\text{if $\dim(x)=\dim(w)$ and $|{\left\langle w,x\right\rangle}|\geq\gamma$,}\\ \star&\text{otherwise}\\ \end{cases}.$
By the aforementioned results of [MP43, Ros58, LNUZ20, KKL+24], the class $\mathcal{H}^{\infty}_{\gamma}$ is pure DP-learnable, shared-randomness replicable, and satisfies $\operatorname{Ldim}(\mathcal{H}^{\infty}_{\gamma})<\gamma^{-2}$. However, by Theorem 2.1, we have $\operatorname{LR}(\mathcal{H}^{\infty}_{\gamma})=\infty$. This establishes the claim. ∎
Disambiguations of large-margin half-spaces.
A disambiguation of a partial concept class $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is a total concept class $\overline{\mathcal{C}}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$ such that for every $c\in\mathcal{C}$ and every finite $S\subseteq c^{-1}(\left\{\pm 1\right\})$, there exists an $\bar{c}\in\overline{\mathcal{C}}$ that is consistent with $c$ on $S$. Intuitively, this corresponds to resolving each $\star$ with 1-1 or +1+1, although this intuition is not completely rigorous in the infinite case.
Disambiguation cannot decrease the list-replicability number. At the same time, it converts a partial class into a total class, where the highly nontrivial results of [BLM20, ABL+22, GGKM21] show that list replicability is bounded in terms of the Littlestone dimension. This principle underlies our results on disambiguations of the large-margin half-space problem and its discrete analogue $\operatorname{GHD}_{\gamma}$. The following theorem resolves an open question of [AHHM21, Question 4] who asked whether every disambiguation of $\mathcal{H}_{\gamma}^{d}$ satisfies $\operatorname{Ldim}=\omega(1)$.
@@ 2.1 Applications
As was noted in [CMY23], it is implicitly  [^3] See [GGKM21, Lemma 5.5] and the definitions of $k_{t}$ and $k^{\prime}$, where $k^{\prime}$ depends on $n_{0}$, $d_{L}$, $\alpha_{\Delta}$, $C_{0}$, and $\eta^{2}$ as specified in [GGKM21, Algorithm 2]. proved in [GGKM21] that for every class $\overline{\mathcal{H}}$,
$\operatorname{LR}_{\epsilon}(\overline{\mathcal{H}})\leq 2^{O_{\epsilon}\left(\operatorname{Ldim}(\overline{\mathcal{H}})^{2}\right)}.$
Recall that by Theorem 2.1, for every $\epsilon\in(0,1/2)$, $\operatorname{LR}_{\epsilon}(\overline{\mathcal{H}})\geq\frac{d}{2}+1$. Combining the two inequalities for a fixed $\epsilon\in(0,1/2)$ concludes the claim. ∎
[AHHM21] used a sophisticated construction, building on Göös’ breakthrough refutation of the Alon–Saks–Seymour conjecture [Göö15], to exhibit partial concept classes with $\operatorname{VCdim}=2$ whose disambiguations satisfy $\operatorname{Ldim}=\omega(1)$. Subsequently, [CHHH23] employed a similar approach to construct partial classes with $\operatorname{Ldim}=2$ whose disambiguations again satisfy $\operatorname{Ldim}=\omega(1)$. Theorem 2.3 provides a much more natural example of a class exhibiting this phenomenon.
Moreover, [CHHH23, Question 4.1] asked whether such a separation can already occur for partial classes of Littlestone dimension $1$. The following corollary of Theorem 2.3 answers this question in the affirmative.
@@
Let $\gamma\in(1/\sqrt{2},1)$. Then $\operatorname{Ldim}(\mathcal{H}_{\gamma}^{d})=1$, while any disambiguation of $\mathcal{H}_{\gamma}^{d}$ has Littlestone dimension $\Omega(\sqrt{\log d})$.
@@
We have $1\leq\operatorname{Ldim}(\mathcal{H}_{\gamma}^{d})\leq 1/\gamma^{2}<2$. An application of Theorem 2.3 completes the proof. ∎
Disambiguations of gap Hamming distance.
In complexity theory, separations between complexity measures are often easier to demonstrate for partial functions, and disambiguations of important partial functions are studied as a way to extend such results to the total setting. For fixed $\gamma\in(0,1)$ the partial $\operatorname{GHD}_{\gamma}$ is known to separate constant cost randomized communication complexity from several other important complexity measures [HHH23, CLV19, HHM23, Son14]. Motivated by this, researchers have asked whether $\operatorname{GHD}_{\gamma}$ admits a disambiguation with constant cost randomized communication complexity [FGHH25]. Our next theorem gives a negative answer to this question.
Theorem 2.5.
$\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n}),$ (3)
and has public-coin randomized communication complexity $\Omega(\log\log n)$.
To lower bound the Littlestone dimension, we use an embedding due to [HHM23] that allows us to disambiguate $\mathcal{H}_{\gamma}^{d}$ using a disambiguation of the Gap Hamming Distance problem in dimension $O(d)$. The key here is that the embedding will maintain the lower bound on Littlestone dimension. Equation 3 then follows from Theorem 2.3. The bound on the communication complexity then follows as a corollary, using the known relationship between Littlestone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity. See Section 4 for the proof.
Pseudo-determinism vs randomness in communication.
A pseudo-deterministic algorithm is a randomized algorithm that, when executed multiple times on the same input, produces the same output with high probability. This notion was introduced by Gat and Goldwasser [GG11] and has since been extensively investigated across a variety of computational models, including learning algorithms, communication protocols, decision tree algorithms, sequential and parallel algorithms, average-case and approximation algorithms, interactive proofs, low-space algorithms, and streaming algorithms (see [GGMW20, GIPS21] and the references therein). A central question in this line of research is to understand to what extent pseudo-determinism can be separated from general randomized computation.
In communication complexity, a search problem is specified by a relation $\mathcal{R}\subseteq\mathcal{X}\times\mathcal{Y}\times\mathcal{Z}$, where Alice and Bob receive $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ respectively, and must output $z\in\mathcal{Z}$ such that $(x,y,z)\in\mathcal{R}$ while minimizing communication. In the public-coin randomized model, the players have access to a shared random string $\bm{r}$, and a protocol $\pi$ must satisfy
$\Pr_{\bm{r}}[(x,y,\pi(\bm{r},x,y))\in\mathcal{R}]\geq\frac{2}{3}\ \ \forall(x,y)\in\mathcal{X}\times\mathcal{Y}.$
Such a protocol is called pseudo-deterministic if there exists a function $f:\mathcal{X}\times\mathcal{Y}\to\mathcal{Z}$ (with $(x,y,f(x,y))\in\mathcal{R}$ for all $(x,y)$) such that
$\Pr_{\bm{r}}[\pi(\bm{r},x,y)=f(x,y)]\geq\frac{2}{3}\ \ \forall(x,y)\in\mathcal{X}\times\mathcal{Y}.$
A well-known candidate for separating pseudo-determinism from randomized communication is the approximate Hamming distance problem
$(x,y,t)\in\operatorname{AHD}_{n}\ \Longleftrightarrow\ |d_{H}(x,y)-t|<\tfrac{n}{3},$
where $d_{H}(x,y)$ is the Hamming distance between $x,y\in\{\pm 1\}^{n}$. In the public-coin randomized model, Alice and Bob can solve $\operatorname{AHD}_{n}$ with only $O(1)$ bits of communication by sampling $O(1)$ coordinates uniformly at random and exchanging the corresponding entries to estimate $d_{H}(x,y)$.
By contrast, it is widely believed that the pseudo-deterministic communication complexity of $\operatorname{AHD}_{n}$ is large. The following theorem establishes the first super-constant lower bound on this problem, thereby yielding the first $O(1)$ versus $\omega(1)$ separation between randomized and pseudo-deterministic communication complexities.
Theorem 2.6.
For any $\epsilon<\frac{1}{2}$, the pseudo-deterministic communication complexity of $\operatorname{AHD}_{n}$ is $\Omega(\log\log(n))$.
Proof.
Suppose there is a pseudo-deterministic protocol for $\operatorname{AHD}_{n}$ of cost $k$, with corresponding function $f:\{\pm 1\}^{n}\times\{\pm 1\}^{n}\to\{0,1,\ldots,n\}$. Define
$F(x,y)\coloneqq\begin{cases}1&f(x,y)\geq\frac{n}{2}\\ -1&f(x,y)<\frac{n}{2}\end{cases},$
and note that the public-coin randomized communication complexity of $F$ is at most $k$.
Since $f(x,y)$ is always within $n/3$ of the true Hamming distance $d_{H}(x,y)$, the function $F$ disambiguates the Gap Hamming Distance problem $\operatorname{GHD}_{0.1}^{n}$, and by Theorem 2.3, its randomized communication complexity is $\Omega(\log\log n)$. ∎
List replicability of finite hyperplane arrangements.
@@
Define the finitary list replicability number of a concept class $\mathcal{C}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$ as
$\widetilde{\operatorname{LR}}(\mathcal{C})\coloneqq\sup_{\text{finite }S\subseteq\mathcal{X}}\operatorname{LR}(\mathcal{C}|_{S}).$
As an example, consider the class $\mathcal{H}^{2}$ of homogeneous halfspaces in $\mathbb{R}^{2}$. While $\operatorname{LR}(\mathcal{H}^{2})=\infty$, Chase, Moran, and Yehudayoff [CMY23, Theorems 5 and 13] proved that for any finite set of points $S\subseteq\mathbb{S}^{1}$, we have $\operatorname{LR}(\mathcal{H}^{2}_{0}|_{S})\leq 2$. This establishes the following striking gap:
@@
They further asked whether a similar bound holds for $\widetilde{\operatorname{LR}}(\mathcal{H}^{3})$ and, more generally, in higher dimensions. The next theorem resolves their question affirmatively.
@@
The upper bound is an easy consequence of the upper bound of our main theorem (Theorem 2.1). Indeed, any finite set of points $S\subset\mathbb{S}^{d-1}$ and hypotheses $H\subset\mathcal{H}^{d}|_{S}$ defined by unit vectors $W\subset\mathbb{S}^{d-1}$ is a sub-concept class of $\mathcal{H}^{d}_{\gamma}$, where $\gamma\coloneqq\min_{x\in S,w\in W}|{\left\langle w,x\right\rangle}|$. The claim now follows from our upper bound from Theorem 2.1.
For the lower bound, we use a result of Chase, Moran, and Yehudayoff [CMY23, Theorem 3] stating that for every concept class $\mathcal{C}$,
$\operatorname{LR}(\mathcal{C})\geq\operatorname{VCdim}(\mathcal{C}).$
The result follows as $\operatorname{VCdim}(\mathcal{H}^{d})=d$, and hence, $\mathcal{H}^{d}$ has a finite subclass of VC dimension $d$. ∎
Theorem 2.7 reveals a connection between list replicability and one of the most fundamental parameters in learning theory called sign-rank. Geometrically, sign-rank is the smallest dimension in which the matrix is realized as points and homogeneous half-spaces.
@@
The sign-rank of a partial class $\mathcal{C}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, denoted by $\operatorname{sign-rank}(\mathcal{C})$, is the smallest $d$ such that there exist vectors $u_{c},v_{x}\in\mathbb{R}^{d}$ for all pairs $c\in\mathcal{C},x\in\mathcal{X}$ such that $c(x)=\operatorname{sgn}({\left\langle u_{c},v_{x}\right\rangle})$ whenever $c(x)\neq\star$.
Combining Theorem 2.7 with the VC-dimension lower bound of [CMY23, Theorem 3] yields the following general bounds on the finitary list replicability number.
@@
For every partial class $\mathcal{C}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, we have
$\operatorname{VCdim}(\mathcal{C})\leq\widetilde{\operatorname{LR}}(\mathcal{C})\leq\operatorname{sign-rank}(\mathcal{C}).$